1. What is the yield to maturity for a $1,000 par, 15 year, 8% coupon bond with annual payments, callable in 3 years for $1,050 that sells for $950?

A. 8.61%

B. 11.55%

C. 3.22%

D. 3.77%

2. What is the yield to call for a $1,000 par, 15 year, 8% coupon bond with annual payments, callable in 3 years for $1,050 that sells for $950?

A. 11.55%

B. 3.77%

C. 8.61%

D. 3.22%

3. What is the yield to maturity for a $1,000 par, 5 year, 6% coupon bond with annual payments, callable in 2 years for $1,050 that sells for $1,100?

A. 3.22%

B. 11.55%

C. 8.61%

D. 3.77%

4. What is the yield to call for a $1,000 par, 5 year, 6% coupon bond with annual payments, callable in 2 years for $1,050 that sells for $1,100?

A. 11.55%

B. 3.22%

C. 3.77%

D. 8.61%

Thank you in advance!!
Answer #1

1
=RATE(15,8%*1000,-950,1000)=8.60593710077676%

2
=RATE(3,8%*1000,-950,1050)=11.5539156975475%

3
=RATE(5,6%*1000,-1100,1000)=3.76832791803776%

4
=RATE(2,6%*1000,-1100,1050)=3.21780515443187%